OpenStudy (anonymous):

Simplify : (3x-1)/(3x-1)^1/2

OpenStudy (anonymous):

\[\frac{a^b}{a^c} = a^{b-c}\]

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

as polpak gave u the law to use (3x-1) ----------- = (3x - 1)^1-1/2 = (3x - 1)^ 1/2 (3x-1)^1/2

OpenStudy (anonymous):

ok thanx